Awolflover1 (awolflover1):

Which shows three ordered pairs that are all solutions of the equation y = x + 7? A. (0, 7), (2, 9), and (3, 11) B. (0, 7), (2, 9), and (3, 10) C. (1, 7), (2, 9), and (3, 10) D. (1, 8), (2, 9), and (3, 11)

OpenStudy (aum):

Substitute the x and y values into the equation y = x + 7 and see if satisfies the equation.

OpenStudy (anonymous):

B. works by inspection.

OpenStudy (aum):

For example, is (0,7) a solution to y = x + 7? put x = 0 and see if you get y = 7. y = 0 + 7 = 7. So (0.7) is a solution. But test the other two in A to see if all three are solutions. Then test the others.